# Unveiling Attacks Ahead of Exploit: adversarial Contract Detection Based on DeFi Attack Features

### 14 Jan 2024 v1

@@ Abstract
We build a dataset consists of features extracted from 304 adversarial contracts and 13,000 benign contracts. Based on this dataset, we evaluate different classifiers, the results of which show that our method for identifying DeFi adversarial contracts performs exceptionally well. For example, the F1-Score for LightGBM-based classifier is 0.9434, with a remarkably low false positive rate of only 0.12%.
1 Introduction
Decentralized Finance (DeFi) has gained significant prominence and is evolving rapidly within the blockchain ecosystem in recent years, giving rise to a diverse array of DeFi applications such as decentralized exchanges (DEXs) [25, 43] and lending platforms [7, 49]. The total value locked (TVL) within these DeFi protocols exceeded \$50 billion in early 2024 [36]. Among various blockchain networks, Ethereum [56] and Ethereum Virtual Machine (EVM)-compatible chains [28, 26, 44, 45] have a thriving DeFi ecosystem, playing a pivotal role in the landscape.
At the heart of DeFi are smart contracts, which are self-executing programs deployed on blockchains, enabling trustless and transparent financial interactions without the need for intermediaries. They could be invoked by users through transactions and autonomously execute predefined actions. However, like traditional software, they are fundamentally code-based programs, making them susceptible to potential security risks. Indeed, DeFi incidents resulting from various smart contract vulnerabilities have caused over 3 billion USD in losses [62, 18, 23]. Figure 1 depicts two common attack patterns:
@@ 1 Introduction
Figure 1: Two common patterns for DeFi attacks: (a) directly interacts with the victim contracts (b) interacts through adversarial contracts.
@@
    Attacking directly. The attacker directly invokes the functions of the victim contracts [11, 13]. However, in the presence of complex attacking logic involving multiple transactions, there is a potential risk of half-way failure. This situation arises when part of the attack is executed but the rest cannot proceed (e.g., due to transaction conflictions).
@@
    Attacking through adversarial contracts. The attacker first deploys an adversarial contract that contains specific attack scripts for the victim contracts, and then initiates the attack by calling the entry function of the adversarial contract[10, 12, 8, 24]. This pattern allows the attacker to execute complex attacking logic in a single transaction, hence avoid half-way failure due to the atomic nature of a transaction. Most DeFi attacks follow this pattern.
Existing defense mechanisms leverage heuristic [60, 63, 46, 58] or machine learning [37, 54] algorithms to detect adversarial transactions, but they face significant challenges in detecting private adversarial transactions. Namely, attackers can send adversarial transactions directly to miners, evading visibility within the blockchain network and effectively bypassing the detection. Based on our analysis of historical DeFi incidents, we found that the proportion of adversarial transactions sent using privacy services has exhibited a significant upward trend, rising from 33% in the first half of 2022 to a notable 71% in the latter half of 2023.
@@
We first collected 304 adversarial contracts used in DeFi incidents between April 2020 and December 2023 from both the Ethereum and Binance Smart Chain (BSC). We also collected the top 13,000 contracts on Ethereum from June 2022 to June 2023, to serve as our benign contract samples. These data are made publicly available to facilitate future research in this domain.
We observe that most adversarial contracts follow a similar pattern, e.g., anonymous fund source, closed-source, frequent token-related function calls. Based on this observation, we extract six categories of features from each contract, such as attributes of contract creator, implemented functions in contract, and contract types. We then build classifiers using multiple supervised machine learning algorithms and evaluate them comprehensively. The results show that our method for identifying DeFi adversarial contracts performs exceptionally well. For example, the F1-Score for LightGBM-based classifier is 0.9434, with a remarkably low false positive rate of only 0.12%.
@@ 2 Background
2.1 Ethereum
Ethereum[56] is an open-source, decentralized blockchain platform that introduces EVM to support smart contracts and has spawned a series of EVM-compatible chains.
Account. Accounts are entities that can hold Ether and send transactions. Ethereum has two main types of accounts: External Owned Account (EOA) and Contract Account (CA). EOAs are controlled by users’ private keys. Both types of accounts have their address, balance, and nonce.
Transcation. Ethereum is often regarded as a transaction-driven state machine, where transactions are initiated by EOA, signed, and ultimately update the Ethereum network’s state. Sending transactions require Ether as transaction fees. Transactions are typically used for transferring Ether, invoking contracts, or creating contracts. In the case of contract deployment, the to field is 0. During the execution of a normal transaction, interacted contracts may trigger additional message calls, commonly known as internal transactions.
Smart Contract. From a high-level perspective, smart contracts can be viewed as instantiated objects stored on the blockchain and executed within the EVM. Due to the nature of blockchain, once deployed, the code within a smart contract cannot be changed. Smart contracts are typically written in high-level programming languages such as Solidity [35]. The source code is compiled to bytecode, which can be executed within the EVM. The EVM is a stack-based virtual machine that supports a set of Turing-complete instructions. Bytecode mainly consists of two parts, creation code and runtime code. Creation code is used for initialization (not deployed on the blockchain) and is executed only once. Runtime code is the primary code segment executed when the contract receives a message call.
2.2 DeFi
@@
The tokens. In addition to Ether, there are numerous cryptocurrencies created by smart contracts on Ethereum, known as tokens. Tokens are the foundation of the DeFi ecosystem, supporting transactions and activities within DeFi. Tokens come in two main types: fungible tokens and non-fungible (NFT) tokens, which widely follow the ERC-20 [1] and ERC-721 [5] standards respectively. A token contract typically must implement a set of functions and events specified in these standards. Additionally, Ether can also be called native token.
@@
Decentralized Exchange (DEX). Similar to traditional markets, the cryptocurrency market depends on exchanges. But unlike centralized exchanges (CEX), DEX doesn’t require users to deposit digital assets into the exchange. Instead, it conducts asset transactions directly on the blockchain through smart contracts, ensuring transparency, security, and decentralization.
@@
Flashloan. Flashloan is a unique form of cryptocurrency lending, allowing users to borrow and repay funds within a single transaction. Many DeFi protocols [7, 2, 25] currently provide flashloan function, offering everyone the opportunity to temporarily possess a large number of tokens with only a small fee. However, flashloan also facilitates numerous attacks[30, 48, 55].
Figure 2: The proportion of private adversarial transactions in 2022 and 2023.
2.3 Miner Extractable Value
@@
Traditionally, miners determine the order of transactions in a block based on transaction fees and ending transactions in blockchain networks are transparent and publicly accessible. Therefore, users have the ability to influence the placement of their transactions by adjusting the transaction fees, allowing for front-running or back-running. The value derived from manipulating the order of transactions within a block, beyond the standard block rewards and transaction fees, is referred to as Miner Extractable Value (MEV) [34]. It has been proven to have a significant impact on the security of blockchain networks [47].
@@
In response to MEV, researchers have proposed various countermeasures [59]. The MEV-auction solution introduced by Flashbots [20] gained popularity and aimed to facilitate efficient, decentralized, and transparent MEV extraction. Users could utilize their private service [42] to send transactions directly to miners, evading visibility within the blockchain network. However, private service providers like Flashbots on Ethereum and 48 club [14] on BSC, are also utilized by attackers to send private adversarial transactions, aiming bypass various detection effectively.
@@
Figure 3: Kernel density estimate of interaction count for contracts created from June 2022 to June 2023
3 Data Collection
@@
Figure 4: The workflow of our methodology for the adversarial contract detection.
@@
3.1 Adversarial Contract
Data Sources. To comprehensively collect adversarial contracts, we gathered information on DeFi incidents from various publicly available sources, such as repositories on GitHub that document historic DeFi attacks [18, 23], attack alerts posted by security companies [51, 15, 16]. Subsequently, we undertake a meticulous manual review of these incidents to identify the adversarial contracts utilized by attackers. Additionally, We also conduct a thorough scrutiny of the transaction history of EOAs involved in these incidents, including the contract creators, callers, and associated transfer accounts, in order to identify any potentially overlooked attack incidents.
@@ 3.1 Adversarial Contract
Scope and Exclusions. Front-running bots often deploy a generic contract that interprets the input data of transactions, akin to a virtual machine, to execute specific logic of the target transaction to frontrun. These contracts do not contain actual vulnerability exploitation logic and therefore are not within the scope of our detection efforts.
Furthermore, we observed that while attack incidents occurred on various blockchains such as Ethereum, BSC, Polygon, and Avalanche, Ethereum and BSC accounted for the vast majority of the data we collected. To streamline further work, we focused on attack incidents from Ethereum and BSC.
It is essential to mention that some attack incidents did not involve contract deployment. For instance, attacks resulting from private key leaks or attackers directly invokes the functions of the victim contracts using EOA. Clearly, these types of attacks fall outside the scope of our detection. Additionally, we only considered contracts containing core attacking logic. Other contracts, such as those used solely for concealing profits to avoid detection or fake token contract to facilitate attacks, were not included in our analysis.
Ultimately, we obtained a total of 304 contracts related to DeFi attacks on Ethereum and BSC from April 2020 to December 2023, with 115 on Ethereum and 189 on BSC, resulting in a cumulative loss estimated to exceed 100 million USD.
Private transactions analysis. By obtaining the labels of transactions related to DeFi incidents through Etherscan, we have the capability to scrutinize the utilization of private services in these attacks. Our analysis revealed that in the DeFi incidents on Ethereum, over 50% of the attacks involved the use of private transactions. Notably, within this specific subset, nearly all of the attacks employed private services when initiating adversarial transactions. Moreover, as depicted in Figure 2, the percentage of adversarial transactions sent using privacy services displayed a remarkable upward trend, surging from 33% in the first half of 2022 to a notable 71% in the latter half of 2023.
3.2 Benign Contract
@@
Furthermore, we conducted interaction count statistics for these contracts using the API provided by Etherscan[19]. The interaction encompass both normal transactions and internal transactions. Due to API constraints, we could retrieve a maximum of 10,000 transactions for each transaction type. The results are depicted in Figure 3. In the entirety of contracts within the specified year, the majority exhibited low interaction counts, with over a quarter of contracts having only a single transaction, which typically corresponds to the contract deployment itself, without any subsequent interactions.
To constitute our benign contract samples, we selected the top 13,000 contracts with the highest interaction counts. The minimum number of interactions among them is close to 1000. We believe that contracts with a high number of interactions are less likely to be associated with attacks, and they exhibit a certain degree of representativeness.
4 Details
@@
4.1 Feature Selection
Table 2 shows a subset of features we selected to classify contracts based on DeFi attack behaviors.
4.1.1 Creator Attribute
Creator Attribute reflects the features of the contract creator. The nonce indicates the number of transactions sent from the contract creator before the deployment of the contract. A higher nonce may suggest a more active EOA address, while attackers often use a brand-new address for their attacks. We also obtain information about the contract creator’s fund source. Typically, attackers acquire funds anonymously, such as through mixers (e.g, Tornado Cash), whereas regular users are more likely to source their funds from KYC-required centralized exchanges.
4.1.2 Transaction Data
Transaction Data includes information related to the transaction that created the contract. In cases where the contract’s constructor is payable, the sender has the capability to directly transfer native tokens, such as ETH on Ethereum, to the contract address during the contract’s creation. Typically, attackers do not need to transfer native tokens to the adversarial contract.
Within the context of the contract deployment, the input data represents the bytecode. The gas consumption depends on the specific content of the data. Consequently, the length of the input data and the corresponding gas consumption could serve as indicators of the complexity of the contract.
We do not select gas price as a feature because it fluctuates dynamically with network congestion. What may be considered a high gas price at one time may not be so at another. Additionally, attackers usually do not increase the gas price to prevent contract deployment from being frontrun.
Table 1: Flashloan Callback and Token-related Function
Flashloan Callback ERC-20 DEX
uniswapV2Call
totalSupply
uniswapV3FlashCallback
balanceOf
uniswapV3SwapCallback
transfer
swap
pancakeCall
allowance
sync
pancakeV3FlashCallback
approve
skim
DPPFlashLoanCall
transferFrom
getReserves
DVMFlashLoanCall
mint
token0
executeOperation
burn
token1
receiveFlashLoan
withdraw
onFlashLoan
deposit
BiswapCall
Table 2: Selected Features
Feature Type Name Description
@@ 4.1.2 Transaction Data
4.1.3 Verification Status

### 02 Feb 2024 v2

# LookAhead: Preventing DeFi Attacks via Unveiling Adversarial Contracts

@@ Abstract.
We build a dataset consists of features extracted from 269 adversarial contracts and 13,000 benign contracts. Based on this dataset, we evaluate different classifiers, the results of which show that our method for identifying DeFi adversarial contracts performs exceptionally well. For example, the F1-Score for LightGBM-based classifier is 0.9541, with a remarkably low false positive rate of only 0.15%.
copyright: noneconference: ; October 14–18, 2024; Salt Lake City, USA
1. Introduction
Decentralized Finance (DeFi) has gained significant prominence and is evolving rapidly within the blockchain ecosystem in recent years, giving rise to a diverse array of DeFi applications such as decentralized exchanges (DEXs) (UniswapV3, 3, 44), lending platforms (AAVE, 2, 52), and derivatives (dydx, 25). The total value locked (TVL) within these DeFi protocols exceeded \$50 billion in early 2024 (DefiLlama, 62). Among various blockchain networks, Ethereum (wood2014ethereum, 68) and Ethereum Virtual Machine (EVM)-compatible chains (Arbitrum, 4, 10, 45, 47) have a thriving DeFi ecosystem, playing a pivotal role in the landscape.
@@ 1. Introduction
Figure 1. Two common patterns for DeFi attacks: (a) directly interacts with the victim contracts (b) interacts through adversarial contracts.
At the heart of DeFi are smart contracts, which are self-executing programs deployed on blockchains, enabling trustless and transparent financial interactions without the need for intermediaries. They could be invoked by users through transactions and autonomously execute predefined actions. However, like traditional software, they are fundamentally code-based programs, making them susceptible to potential security risks. Indeed, DeFi incidents resulting from various smart contract vulnerabilities have caused over 3 billion USD in losses (DeFiHackLabs, 24, 42, 75). Figure 1 depicts two common attack patterns:
@@
    Attacking directly. The attacker directly invokes the functions of the victim contracts (TreasureDAOHack, 6, 59). However, in the presence of complex attacking logic involving multiple transactions, there is a potential risk of half-way failure. This situation arises when part of the attack is executed but the rest cannot proceed (e.g., due to transaction conflictions).
@@
    Attacking through adversarial contracts. The attacker first deploys an adversarial contract that contains specific attack scripts for the victim contracts, and then initiates the attack by calling the entry function of the adversarial contract (BeanstalkHack, 5, 11, 46, 58). This pattern allows the attacker to execute complex attacking logic in a single transaction, hence avoid half-way failure due to the atomic nature of a transaction. Most DeFi attacks follow this pattern.
Existing defense mechanisms leverage heuristic (qin2023towards, 49, 70, 72, 76) or machine learning (gai2023blockchain, 37, 64) algorithms to detect adversarial transactions, but they face significant challenges in detecting private adversarial transactions. Namely, attackers can send adversarial transactions directly to miners, evading visibility within the blockchain network and effectively bypassing the detection. Based on our analysis of historical DeFi incidents, we found that the proportion of adversarial transactions sent using privacy services has exhibited a significant upward trend, rising from 33% in the first half of 2022 to a notable 63% in the latter half of 2023.
@@
We first collected 269 adversarial contracts used in DeFi incidents between April 2020 and December 2023 from both the Ethereum and Binance Smart Chain (BSC). We also collected the top 13,000 contracts on Ethereum from June 2022 to June 2023, to serve as our benign contract samples. These data are made publicly available to facilitate future research in this domain.
We observe that most adversarial contracts follow a similar pattern, e.g., anonymous fund source, closed-source, frequent token-related function calls. Based on this observation, we extract six categories of features from each contract, such as attributes of contract creator, implemented functions in contract, and contract types. We then build classifiers using multiple supervised machine learning algorithms and evaluate them comprehensively. The results show that our method for identifying DeFi adversarial contracts performs exceptionally well. For example, the F1-Score for LightGBM-based classifier is 0.9541, with a remarkably low false positive rate of only 0.15%.
@@ 2. Background
2.1. Ethereum
Blockchain is a distributed ledger built on a peer-to-peer (P2P) network, maintained through specific consensus mechanisms to ensure the consistency of the ledger. Ethereum (wood2014ethereum, 68) is one of the most renowned blockchain platforms. It introduces Ethereum Virtual Machine (EVM) to support smart contracts and has spawned a series of EVM-compatible chains.
Account. Accounts are entities that can hold Ether and send transactions. Ethereum has two types of accounts: External Owned Accounts (EOAs) and Contract Accounts (CAs). EOAs are controlled by users’ private keys, while CAs contain corresponding code and storage. Both types of accounts have their address, balance, and nonce. The address of an EOA is derived from the hash of the public key, whereas the address of a CA is usually related to its creator.
Transcation. Ethereum is often regarded as a transaction-driven state machine, where transactions are initiated by EOA, signed, and ultimately update the Ethereum network’s state. After a transaction is sent, it’s first placed in the mempool, a transparent and publicly accessible pool filled with pending transactions, waiting for the block producer to select and finalize them. Transactions are typically used for transferring Ether, invoking contracts, or creating contracts. In the case of contract deployment, the to field of the transaction is 0. During the execution of a normal transaction, interacted contracts may trigger additional message calls, commonly known as internal transactions.
Smart Contract. From a high-level perspective, smart contracts can be viewed as instantiated objects stored on the blockchain and executed within the EVM. Due to the nature of blockchain, once deployed, the code within a smart contract cannot be changed. Smart contracts are typically written in high-level programming languages such as Solidity (dannen2017introducing, 22). The source code is compiled to bytecode, which can be executed within the EVM. The EVM is a stack-based virtual machine that supports a set of Turing-complete instructions. Bytecode mainly consists of two parts, creation code and runtime code. Creation code is used for initialization (not deployed on the blockchain) and is executed only once. Runtime code is the primary code segment executed when the contract receives a message call.
2.2. DeFi
@@
The tokens. There are numerous cryptocurrencies created by smart contracts on Ethereum, known as tokens. These tokens serve as the cornerstone of the DeFi ecosystem, facilitating the diversified development of applications and services within DeFi. Tokens come in two main types: fungible tokens and non-fungible tokens (NFT), which widely follow the ERC-20 (ERC20, 28) and ERC-721 (ERC721, 30) standards respectively. A token contract typically must implement a set of functions and events specified in these standards. Furthermore, Ether, which is used to reward contribution in the consensus, is also referred to as the native token.
@@
Decentralized Exchange (DEX). Similar to traditional markets, the cryptocurrency market depends on exchanges. But unlike centralized exchanges (CEX), DEX doesn’t require users to deposit digital assets into the exchange. Instead, it conducts asset transactions directly on the blockchain through smart contracts, ensuring transparency, security, and decentralization. The most prevalent DEXs (UniswapV3, 3, 44) utilize the Automated Market Maker (AMM) model, which relies on mathematical formulas to dynamically adjust the prices of digital assets, diverging from the traditional order book approach used in CEXs. In this model, the core contracts known as liquidity pools are funded by entities called liquidity providers, allowing users to continuously exchange assets.
@@
Flashloan. Flashloan is a unique form of cryptocurrency lending that leverages the atomic nature of the transaction, where a transaction is either fully executed or not executed at all, allowing users to borrow and repay cryptocurrency in one transaction. To utilize the flashloan, specific callback functions must be implemented in the borrowing smart contracts. Many DeFi protocols (AAVE, 2, 3, 25) currently provide flashloan function, offering everyone the opportunity to temporarily possess a large number of tokens with only a small fee. However, flashloan also facilitates numerous attacks (cao2021flashot, 13, 51, 65).
Figure 2. The proportion of private adversarial transactions in 2022 and 2023.
2.3. Miner Extractable Value
@@
Traditionally, miner determine the order of transactions in a block based on transaction fees. Since the mempool is publicly accessible, users have the ability to influence the placement of their transactions by adjusting the transaction fees, allowing for front-running (eskandari2020sok, 32, 60) or back-running. The value derived from manipulating the order of transactions within a block, beyond the standard block rewards and transaction fees, is referred to as Miner Extractable Value (MEV) (daian2020flash, 21). It has been proven to have a significant impact on the security of blockchain networks (qin2022quantifying, 50).
Front-running. Front-running is a fundamental means used to extract MEV, typically employed in arbitrage, liquidation, and sandwich attack. Front-runners usually deploy a contract on the blockchain, akin to a virtual machine, capable of interpreting the input data of transactions. When the front-running bot detects a victim transaction that meets specific conditions, it constructs input data to execute the specific logic for profit. On the other hand, front-running has also been considered as a means in the intrusion prevention systems (qin2023blockchain, 48, 73), attempting to frontrun adversarial transactions to prevent the attacks from occurring.
@@
Private Service. In response to MEV, researchers have proposed various countermeasures (yang2022sok, 71). The MEV-auction solution introduced by Flashbots (Flashbots, 34) gained popularity (weintraub2022flash, 67). Users could utilize their service (lyu2022empirical, 43) to send private transactions directly to miners, evading exposure in the mempool. However, private service providers like Flashbots on Ethereum and 48 club (bnb48, 1) on BSC, are also utilized by attackers to send private adversarial transactions, aiming bypass various detection effectively.
@@
Figure 3. Kernel density estimate of interaction count for contracts created from June 2022 to June 2023.
3. Data Collection
@@
Figure 4. The workflow of our methodology for the adversarial contract detection.
Category Count (Proportion)
Total Attacks 109
Involving Private 54 (49.54%)
Adversarial Tx Private 52 (96.30% of Private)
Table 1. Analysis of private transactions in the DeFi incidents on Ethereum
@@
3.1. Adversarial Contract
Data Sources. To comprehensively collect adversarial contracts, we gathered information on DeFi incidents from various publicly available sources, such as repositories on GitHub that document historic DeFi attacks (DeFiHackLabs, 24, 42), attack alerts posted by security companies (blocksec, 8, 15, 55). Subsequently, we undertake a meticulous manual review of these incidents to identify the adversarial contracts utilized by attackers. Additionally, We also conduct a thorough scrutiny of the transaction history of EOAs involved in these incidents, including the contract creators, callers, and associated transfer accounts, in order to identify any potentially overlooked attack incidents.
@@ 3.1. Adversarial Contract
Scope and Exclusions. Due to the contracts deployed by front-running bots do not contain actual vulnerability exploitation logic and therefore are not within the scope of our detection efforts. Furthermore, we observed that while attack incidents occurred on various blockchains, Ethereum and BSC accounted for the vast majority of the data we collected. To streamline further work, we focused on attack incidents from Ethereum and BSC. Additionally, we only considered contracts containing the core attacking logic. Other contracts, such as those used solely for concealing profits to avoid detection or fake token contract to facilitate attacks, were not included in our analysis.
Ultimately, we obtained a total of 269 contracts related to DeFi attacks on Ethereum and BSC from April 2020 to December 2023, with 109 on Ethereum and 160 on BSC, resulting in a cumulative loss estimated to exceed 1 billion USD.
Private transactions analysis. Transactions with labels of MEV Transactions or Flashbots in Etherscan indicate that they bypassed the Ethereum mempool and were sent directly to the block producer. By obtaining the labels of transactions related to DeFi incidents through Etherscan, we have the capability to scrutinize the utilization of private transactions in these attacks. As shown in table 1, our analysis revealed that in the DeFi incidents on Ethereum, about 50% of the attacks involved the use of private transactions. Notably, within this specific subset, nearly all of the attacks employed private services when initiating adversarial transactions. Moreover, as depicted in Figure 2, the percentage of adversarial transactions sent using privacy services displayed a remarkable upward trend, surging from 33% in the first half of 2022 to a notable 63% in the latter half of 2023.
3.2. Benign Contract
@@
Furthermore, we conducted interaction count statistics for these contracts using the API provided by Etherscan (etherscan, 33). The interaction encompass both normal transactions and internal transactions. Due to API constraints, we could retrieve a maximum of 10,000 transactions for each transaction type. The results are depicted in Figure 3. In the entirety of contracts within the specified year, the majority exhibited low interaction counts, with over a quarter of contracts having only a single transaction, which typically corresponds to the contract deployment itself, without any subsequent interactions.
To constitute our benign contract samples, we selected the top 13,000 contracts with the highest interaction counts. The minimum number of interactions among them is close to 1,000. We believe that contracts with a high number of interactions are less likely to be associated with attacks, and they exhibit a certain degree of representativeness.
4. Details
@@
4.1. Feature Selection
Feature Category Name Description
@@ 4.1. Feature Selection
Table 2. Selected Features
Table 2 shows the core subset of features we selected to classify contracts based on DeFi attack behaviors, divided into six categories.
4.1.1. Creator Attribute
Figure 5. Distribution of input data length and gas consumption for both adversarial and benign contract samples.
This category encompasses features related to the contract creator. The nonce indicates the number of transactions sent from the contract creator before the deployment of the contract. A higher nonce may suggest a more active EOA address. We also obtain information about the contract creator’s fund source. To conceal their identity, attackers typically transfer initial funds from a mixer (e.g, Tornado Cash (wang2023zero, 66)) or instant exchanges (ChangeNOW, 16, 54) without the need for sign-up or KYC (Know Your Customer), to a new account for subsequent transactions. In contrast, regular users are more likely to source their funds from KYC-required centralized exchanges.
4.1.2. Transaction Data
This category pertains to information concerning the transaction that created the contract. In cases where the contract’s constructor is payable, the creator possesses the ability to directly transfer native tokens, such as ETH on Ethereum, to the contract address during the contract’s creation. The specific amount transferred is denoted by the value field within the transaction.
In the context of the contract deployment, the input data field of a transaction represents the bytecode of the contract. The gas consumption for a transaction is contingent upon the specific content of the bytecode. Consequently, the length of the input data and the corresponding gas consumption could serve as indicators of the complexity of the contract. As depicted in Figure 5, within our contract samples, it is observed that in the case of adversarial contracts, the deployment transactions typically involve less input data, with gas consumption concentrated in the range of 1.0e6 to 1.5e6 wei. Conversely, benign contracts exhibit slightly more input data, accompanied by a more widespread gas consumption, exceeding 2.0e6 in general.
We do not select gas price as a feature because it fluctuates dynamically with network congestion. What may be considered a high gas price at one time may not be so at another. Additionally, attackers usually do not need to increase the gas price to prevent contract deployment from being frontrun.
4.1.3. Verification Status
Verified Unverified
Benign 11594 (89.2%) 1406 (10.8%)
Adversarial 4 (1.5%) 265 (98.5%)
Table 3. Verification status for both benign and adversarial contract samples.